### v2

@@ 2.1 Model
The mechanism designer needs to design a truthful mechanism $\mathcal{M}$ that receives the valuation functions of the agents $v_{1},\ldots,v_{n}$ and returns an allocation $A=(A_{1},\ldots,A_{n})\in\Delta$, where $\Delta$ is the set of all possible allocations and a payment vector $p=(p_{1},\ldots,p_{n})\in\mathbb{R}_{\geq 0}^{n}$, where the mechanism might use randomness, in which case, $A$ and $p$ are random variables returned by the mechanism. The mechanism designer’s goal is to maximize the expected residual surplus, which is $\sum_{i\in\mathcal{N}}v_{i}(A_{i})-p_{i}$.
@@
2.2 The VCG Mechanism with Copies
In this section, we introduce the VCG with copies framework. We then show how to use this framework and devise multiple truthful mechanisms with optimal guarantees for various settings.
Input: An items set $\mathcal{I}$, valuation profile $v_{1},\ldots,v_{n}$, parameters $r\in\mathbb{N},~{}q\in[0,1]$, and a subroutine $B$
@@ 1 Draw $\ell$ uniformly at random from $\{0,\ldots,r\}$
Let =×$\mathcal{I}^{\prime}=\mathcal{I}\times[2^{\ell}]$   // Create $2^{\ell}$ copies of each item
2 Let $g:\mathcal{I}^{\prime}\rightarrow\mathcal{I}$ be the function where $g(S)=\{j\mid\exists k\in[2^{\ell}]:(j,k)\in S\}$
3 Let $v_{i}^{\prime}:\mathcal{I}\rightarrow\mathbb{R}_{\geq 0}$ be the following function where $v_{i}^{\prime}(S)=v_{i}(g(S))$
@@
5 Use the subroutine $B$ on $(A^{\prime},p^{\prime},\ell,q)$ to create an allocation $A$ and a payment vector $p$
  // $B$ guarantees that each agent $i$ receives set $g(A_{i}^{\prime})$ at price $p^{\prime}_{i}$ with probability $\frac{q}{2^{\ell}}$
@@
Mechanism 1 basically creates $2^{\ell}$ copies of all the items, and extends the valuation functions of the agents to the instance with the copies. Then, the mechanism calculates the VCG allocation and price vector for the instance with the copies. Since in the real instance, there is only one copy of each item, in order to create a valid allocation, the mechanism uses the randomized subroutine $B$ to generate an allocation and payments, where each agent receives the equivalent set and pays the same as in the allocation returned by Step 1 with a probability of $\frac{q}{2^{\ell}}$ (and otherwise, the agent receives nothing and pays nothing). We assume that in Step 1, the calculated allocation is non-redundant and all allocated items contribute to the welfare. I.e., $\forall i\in\mathcal{N}$, and $j\in A^{\prime}_{i}$, it holds that $v_{i}^{\prime}(A^{\prime}_{i})>v_{i}(A_{i}^{\prime}\setminus\{j\})$, and in particular, no agent receives two copies of the same item. In the remainder of this section, we show how to implement the subroutine $B$ for two cases, (1) where all agents have unit-demand valuations, and (2) where all agents have multi-unit submodular valuations. We then show that for these cases, the mechanism guarantees a residual surplus of at least $\frac{q}{r+1}$ of the entire social welfare.
We first claim that as long as the subroutine $B$ satisfies the property that each agent $i$ receives set $g(A_{i}^{\prime})$ and pays $p_{i}^{\prime}$ with probability $\frac{q}{2^{\ell}}$ (and receives the empty set and pays nothing otherwise) then Mechanism 1 is truthful in expectation and ex-post individual rational no matter how the subroutine is implemented.
@@
Mechanism 1 is TIE and EPIR.
@@
To show EPIR, we observe that the outcome for agent $i$ is either the emptyset in which case the agent pays nothing which satisfies EPIR, or the agent receives a utility of $v_{i}(A_{i})-p_{i}=v_{i}(g(A_{i}^{\prime}))-p_{i}^{\prime}=v_{i}^{\prime}(A_{i}^{\prime})-p_{i}^{\prime}\geq 0$, where the inequality is by individual rationality of the VCG mechanism.
To show TIE, we observe that since the subroutine does not influence the possible sets that the agent may receive, the payments or the probability of being allocated, it is sufficient to show that reporting $v_{i}$ always maximizes the value of $v_{i}(g(A_{i}^{\prime}))-p_{i}^{\prime}$ for every realization of $\ell$, and any report of $v_{-i}$. Let $\bar{v}_{i}$ be an arbitrary report, and let $(\bar{A}_{i}^{\prime},\bar{p}_{i}^{\prime})$ be the set allocated to agent $i$ and his payment when misreporting $\bar{v}_{i}$ instead of $v_{i}$, calculated in Step 1. Thus,
$v_{i}(g(A_{i}^{\prime}))-p_{i}^{\prime}=v_{i}^{\prime}(A_{i}^{\prime})-p^{\prime}_{i}\geq v_{i}^{\prime}(\bar{A}_{i}^{\prime})-\bar{p}_{i}^{\prime}=v_{i}(g(\bar{A}_{i}^{\prime}))-\bar{p}_{i}^{\prime},$
where the inequality is by the truthfulness of the VCG mechanism, and the two equalities hold since $v_{i}(S)=v_{i}^{\prime}(g(S))$ for every set $S$. This concludes the proof. ∎
@@
In order to implement Mechanism 1 for the case where all agents have unit-demand valuations and $q=1$, we can do the following: For each item independently select a random copy of it (among the $2^{\ell}$ copies). If this copy of the item is allocated to some agent, then allocate the item to this agent and charge accordingly. This is a valid implementation since by our assumption of the non-redundancy of the allocation returned at Step 1 and by the fact that the valuations are unit-demand, so that no agent receives two items. Thus, the allocation $A^{\prime}$ creates a partition of a subset of the agents according to the different items. Thus, the problem can be separated by this partition, which concludes the proof that this is a valid implementation of subroutine $B$.
@@
In order to implement Mechanism 1 for the case where all agents have multi-unit valuations and $q=\frac{1}{2}$, we will need to make an additional assumption regarding the implementation of the VCG mechanism used in Step 1 of the mechanism. The additional assumption is that the items are allocated sequentially where the items are ordered according to $(1,1),\ldots,(n,1),(1,2),\ldots,(n,2),\ldots,(1,2^{\ell}),\ldots,(n,2^{\ell})$, and then the VCG mechanism partitions the items into consecutive intervals $I_{1},\ldots,I_{n},I_{\bot}$, where agent $i$ receives set $I_{i}$, and $I_{\bot}$ is the set of unallocated items (see Figure 1 for an illustration of the implementation of the subroutine). This assumption can be made without loss of generality since the valuations ($v_{1}^{\prime},\ldots,v_{n}^{\prime}$) are basically multi-unit as long as they don’t receive two copies of the same original item, and giving consecutive intervals of items of length at most $n$ guarantees that no two copies of the same item are given to the same agent. Thus, we can reallocate the items in the described way and this is a valid allocation that also maximizes the social welfare when allocating $\mathcal{I}^{\prime}$. In order to implement the subroutine $B$ for $q=\frac{1}{2}$ we do the following: For each copy $t\in[2^{\ell}]$, let $S_{t}=\{i\in\mathcal{N}\mid\emptyset\subset A_{i}^{\prime}\subseteq\mathcal{I}\times\{t\}\}$, and let $S_{t}^{\prime}=\{i\in\mathcal{N}\mid\exists j_{1},j_{2}:~{}(j_{1},t),(j_{2},t+1)\in A_{i}^{\prime}\}$. In other words, $S_{t}$ is the set of agents that received only items from the $t$-th copy (and received at least one item), and $S_{t}^{\prime}$ is the set of agents that received items from both the $t$-th and $(t+1)$-th copies. We then observe that by our assumption that the agents are allocated sequentially, then $S_{t}^{\prime}$ is at most of size 1. Now, in order to implement the subroutine $B$, we choose a uniformly random index $t$ among $[2^{\ell}]$. Then we either allocate to all agents in $S_{t}$ with probability $\frac{1}{2}$, or allocate to the agent in $S_{t}^{\prime}$ (if such exists) with probability $\frac{1}{2}$. Allocation to a set of agents of the form $S_{t}$ is valid since it is composed by only one copy and therefore the sets created by projecting using the function $g$ are disjoint. Allocating to a set of agents of the form $S_{t}^{\prime}$ is valid since it is composed of at most one agent.
@@
[…]
2.3 Analysis of the VCG Mechanism with Copies for Gross-Subtitutes
@@
When Mechanism 1 receives gross-substitute valuations then the expected residual surplus of its output is at least $\frac{q}{r+1}\left(SW(\mathcal{N},\mathcal{I})-\frac{SW(\mathcal{N},2^{r}\cdot\mathcal{I})}{2^{r}}\right).$
@@
We next define the following notation: For an agent $i\in\mathcal{N}$, and a set of items $\mathcal{I}^{\prime}$ which consists of $k$ copies of the original set $\mathcal{I}$, let $p_{i}(\mathcal{N},\mathcal{I}^{\prime})$ be the payment charged to agent $i$ in the VCG mechanism when applied to $\mathcal{N}$ and $\mathcal{I}^{\prime}$ [^3] We require that the payments are consistent with the implementation used in Step 1 of Mechanism 1. . We also denote by $p(\mathcal{N},\mathcal{I}^{\prime})=\sum_{i\in\mathcal{N}}p_{i}(\mathcal{N},\mathcal{I}^{\prime})$. Then, we first prove the following claim:
@@
It holds that $SW(\mathcal{N},2\cdot\bar{\mathcal{I}})-SW(\mathcal{N},\bar{\mathcal{I}})\geq\frac{p(\mathcal{N},2\cdot\bar{\mathcal{I}})}{2}$.
@@
$\leq\sum_{j\in A_{i}(\mathcal{N},2\cdot\bar{\mathcal{I}})}SW(\mathcal{N},2\cdot\bar{\mathcal{I}}+\{j\})-SW(\mathcal{N},2\cdot\bar{\mathcal{I}}),$
@@
$\geq\frac{q}{r+1}\cdot\left(SW(1)\cdot\left(2-\frac{1}{2^{r}}\right)+\sum_{i=1}^{r}(SW(2^{i})-SW(2^{i-1}))\cdot 2\left(\frac{1}{2^{i}}-\frac{1}{2^{r+1}}\right)\right),$ (2)
where the second equality is by telescoping sum; the third equality is by geometric sum; and the inequality is since $SW(\mathcal{N},\mathcal{I})\geq p(\mathcal{N},\mathcal{I})$. On the other hand, the expected payment of the mechanism is:
$\mbox{\bf E}[\sum_{i\in\mathcal{N}}p_{i}]=\frac{q}{r+1}\cdot\sum_{l=0}^{r}\frac{p(2^{l})}{2^{l}}\leq\frac{q}{r+1}\cdot\left(SW(1)+\sum_{l=1}^{r}\frac{SW(2^{l})-SW(2^{l-1})}{2^{l-1}}\right),$ (3)
@@
When Mechanism 1 receives unit-demand valuations then the expected residual surplus of its output is at least $O\left(\log(n)\right)$ approximation to the social welfare.
@@
When Mechanism 1 receives submodular multi-unit valuations then the expected residual surplus of its output is at least $O\left(\log(n)\right)$ approximation to the social welfare.

### v3

@@ 2.1 Model
The mechanism designer needs to design a truthful mechanism $\mathcal{M}$ that receives the valuation functions of the agents $v_{1},\ldots,v_{n}$ and returns an allocation $A=(A_{1},\ldots,A_{n})\in\Delta$, where $\Delta$ is the set of all possible allocations of $\mathcal{I}$ and a payment vector $p=(p_{1},\ldots,p_{n})\in\mathbb{R}_{\geq 0}^{n}$, where the mechanism might use randomness, in which case, $A$ and $p$ are random variables returned by the mechanism. The mechanism designer’s goal is to maximize the expected residual surplus, which is $\sum_{i\in\mathcal{N}}v_{i}(A_{i})-p_{i}$.
@@
2.2 The VCG Mechanism with Copies
In this section, we introduce the VCG with copies framework. We then show how to use this framework and devise multiple truthful mechanisms with optimal guarantees for various settings. We begin by defining the key components of the framework.
Instance with Copies.
Consider an instance with a set $\mathcal{I}$ of items and a set $\mathcal{N}$ of $n$ agents with valuation functions $v_{1},\dots,v_{n}$. We define an instance with $2^{\ell}$ copies of each item in $\mathcal{I}$, yielding an item set =×$\mathcal{I}^{\prime}=\mathcal{I}\times[2^{\ell}]$. We define a projection function $g:2^{\mathcal{I}^{\prime}}\to 2^{\mathcal{I}}$ that maps a subset of copied items back to their original items: $\mathcal{I}$, $g(S)=\{j\mid\exists k\in[2^{\ell}]:(j,k)\in S\}$. We define the extension of $v:2^{\mathcal{I}}\rightarrow\mathbb{R}_{\geq 0}$ to an instance with $2^{\ell}$ copies by $v^{\prime}:2^{\mathcal{I}^{\prime}}\to\mathbb{R}_{\geq 0}$ where $v^{\prime}(S)=v(g(S))$. Throughout the paper, we denote the valuations, sets, and payments of an instance with copies using a prime (i.e., )—while the original instance’s valuations, sets, and payments are denoted without a prime.
VCG mechanism.
@@ 2.2 The VCG Mechanism with Copies
Rounding Scheme.
A key component of the mechanism we devise is a rounding scheme. A rounding scheme takes as input $(A^{\prime},\ell)$, where $A^{\prime}$ is an allocation of an instance with $2^{\ell}$ copies of each item. The rounding scheme (randomly) outputs an allocation $A$ for the original instance (without the copies).
An allocation $A^{\prime}=(A^{\prime}_{1},\ldots,A_{n}^{\prime})$ is non-redundant with respect to class of valuations $\mathcal{C}$ if there is no $S_{i}^{\prime}$ which is a strict subset of $A^{\prime}_{i}$ such that $v(S_{i}^{\prime})=v(A^{\prime}_{i})$ for each $v\in\mathcal{C}$.
We say that a rounding scheme $B$ is $q$-valid with respect to a class $\mathcal{C}$ of valuation functions if for every $\ell\geq 0$ and a non-redundant allocation $A^{\prime}$ of $2^{\ell}$ copies of the items with respect to $\mathcal{C}$, $B$ allocates agent $i$ a (random) set $A_{i}$ such that for every $v\in\mathcal{C}$ it holds that
$\mbox{\bf E}[v(A_{i})]=\frac{q}{2^{\ell}}\cdot v^{\prime}(A^{\prime}_{i}),$ (1)
where $v^{\prime}$ is the extension of $v$ to $2^{\ell}$ copies and the expectation is over the randomness of the rounding scheme.
The main challenge is to show how to devise a valid rounding scheme with constant $q$ for certain classes of valuations. With these components in place, we can now formally define our mechanism.
Input: An item set $\mathcal{I}$, valuation profile $v_{1},\ldots,v_{n}$, parameter $r\in\mathbb{N}$, and a rounding scheme $B$
@@ 1 Draw $\ell$ uniformly at random from $\{0,\ldots,r\}$
Let =×$\mathcal{I}^{\prime}=\mathcal{I}\times[2^{\ell}]$
  // Create $2^{\ell}$ copies of each item
2 Let $v_{i}^{\prime}:2^{\mathcal{I}^{\prime}}\rightarrow\mathbb{R}_{\geq 0}$ be the extension of $v_{i}$ to $2^{\ell}$ copies
@@
4 Use the rounding scheme $B$ on $(A^{\prime},\ell)$ to create an allocation $A$
For every agent $i\in\mathcal{N}$, set $p_{i}=p^{\prime}_{i}\cdot\frac{v_{i}(A_{i})}{v^{\prime}_{i}(A^{\prime}_{i})}$ if $v^{\prime}_{i}(A^{\prime}_{i})>0$ and $p_{i}=0$ otherwise.
@@ 4 Use the rounding scheme $B$ on $(A^{\prime},\ell)$ to create an allocation $A$
Mechanism 1 basically creates $2^{\ell}$ copies of all the items, and extends the valuation functions of the agents to the instance with the copies. Then, the mechanism calculates the VCG allocation and price vector for the instance with the copies. We assume that in Step 1, the calculated allocation is non-redundant with respect to the class of valuations we consider. Since in our actual instance, there is only a single copy of each item, we utilize a $q$-valid rounding scheme $B$ to generate an allocation for the original instance. The rounding scheme guarantees that the expected value that each agent receives is (in expectation) proportional to the value in the allocation returned by Step 1. Lastly, in Step 1, we scale for each agent the payment returned by the VCG by the fraction of value allocated to this agent over the value returned by the VCG.
In the remainder of this section, we show that given a $q$-valid rounding scheme, our mechanism is TIE and EPIR (Claim 1). We then devise $q$-valid rounding schemes for unit-demand valuations with $q=1$ and for multi-unit submodular valuations with $q=1/2$. We then present residual surplus guarantees for the class of gross-substitutes valuations as a function of $q$ and $r$ (Lemma 3). Combining everything, we obtain an $O(\log(n))$-approximation.
The key challenges for establishing guarantees using our framework are: 1) Devising $q$-valid rounding schemes with large values of $q$ which we show is possible for unit-demand (or even additive subject to matroid constraints, see discussion in Section 4) and multi-unit submodular valuations. 2) Comparing the payments and utilities of the outcomes of the VCG mechanism for different amounts of copies. In particular, our key lemma is for the case of gross-substitutes valuations where we show that the additional contribution to the welfare by adding a copy of all items is at least half of the sum of payments of the VCG mechanism on the instance with the additional copies (Claim 6).
We first claim that as long as the rounding scheme $B$ is $q$-valid then Mechanism 1 is truthful in expectation and ex-post individual rational no matter how the rounding scheme is implemented.
@@
If Mechanism 1 receives a $q$-valid rounding scheme $B$ with respect to a class $\mathcal{C}$ of valuation functions, then it is TIE and EPIR with respect to $\mathcal{C}$.
@@
To show EPIR, we observe that since the VCG mechanism is EPIR we have $p^{\prime}_{i}\leq v^{\prime}_{i}(A^{\prime}_{i})$. Using this, if $v^{\prime}_{i}(A^{\prime}_{i})$ is 00, $i$ pays nothing. Otherwise, the payment of agent $i$ for a set $A_{i}$ of items is $p_{i}=p^{\prime}_{i}\cdot\frac{v_{i}(A_{i})}{v^{\prime}_{i}(A^{\prime}_{i})}\leq v_{i}(A_{i})$, meaning that the agent pays no more than his value for the set of items. To show TIE, we fix a value of $\ell$. For a $q$-valid rounding scheme, the expected value of an agent $i$ is $\frac{q}{2^{\ell}}\times v_{i}(g(A^{\prime}_{i}))$, where $A^{\prime}_{i}$ is the allocation to agent $i$ in the VCG mechanism with $2^{\ell}$ copies and $g(A^{\prime}_{i})$ is the projection of those items to the actual instance. By Step 1 and $B$ being a $q$-valid rounding scheme, the expected payment of agent $i$ is the same as in the VCG mechanism with $2^{\ell}$ copies of each item, that is $p^{\prime}_{i}$. Observe that two statements imply that for every agent his expected utility is $\frac{q}{2^{\ell}}$ times his utility in the instance with $2^{\ell}$ copies. Since the VCG mechanism is truthful, reporting $v_{i}$ maximizes the agent’s expected utility. ∎
We next show that for the cases of unit-demand and multi-unit valuations
@@
We construct a $1$-valid rounding scheme $B_{UD}$ for the class of unit-demand valuations, as follows: For each item, independently select a random copy of it (among the $2^{\ell}$ copies). If this copy of the item is allocated to some agent, then allocate the item to this agent and charge accordingly. This is a $q$-valid rounding scheme since any non-redundant allocation returned at Step 1 with respect to the class of unit-demand valuations allocates at most one item per agent. An extension $v^{\prime}:2^{\mathcal{I}^{\prime}}\rightarrow\mathbb{R}_{\geq 0}$ of a unit demand function $v:2^{\mathcal{I}}\rightarrow\mathbb{R}_{\geq 0}$ is unit-demand. Thus, the allocation $A^{\prime}$ creates a partition of a subset of the agents according to the different items. Therefore, allocating each item $j\in\mathcal{I}$ to each of agents a copy of it is allocated to with probability $2\ell$ independently is a $1$-valid rounding scheme for unit-demand valuations.
@@
We construct a $\frac{1}{2}$-valid rounding scheme $B_{MU}$ for the class of multi-unit valuations. Given a non-redundant allocation $A^{\prime}$ with respect to the class of multi-unit valuations, where each item has $2^{\ell}$ copies. We first create $n$ buckets of agents $B_{1},\dots B_{n}$, where at the beginning each bucket contains a single agent. As long as there are two buckets, for which the number of items allocated to the agents within these buckets is at most $m$, we merge them into a single bucket. That is, for each bucket $B$, we maintain the condition: $|\bigcup_{i\in B}A^{\prime}_{i}|\leq m$. Once we stop merging, it must be that for every pair of remaining buckets, the total number of items allocated in every pair of buckets is at least $m+1$. I.e., for any two different buckets $B,B^{\prime}$, it holds that $|\bigcup_{i\in B\cup B^{\prime}}A^{\prime}_{i}|>m$ Since in the instance with the copies we have overall $m\cdot 2^{\ell}$ items, this implies that at the end of the process, at most $\left\lceil 2\frac{m\cdot 2^{\ell}}{m+1}\right\rceil\leq 2\cdot 2^{\ell}$ buckets remain. Now, to implement a $\frac{1}{2}$-valid rounding scheme for multi-unit valuation functions we choose each bucket with probability $\frac{1}{2\cdot 2^{\ell}}$, and allocate only to agents that belong to this bucket the same amount of items as in $A^{\prime}$. (If there are less than $2\cdot 2^{\ell}$ buckets, then with the remaining probability, we allocate nothing.) This allocation is feasible as in each bucket the sum of items allocated to the agents in this bucket is at most $m$. Moreover, each agent is allocated with probability $\frac{1}{2\cdot 2^{\ell}}$, which satisfies Equation (1).
@@
2.3 Analysis of the VCG Mechanism with Copies for Gross-Substitutes
@@
For a subclass $\mathcal{C}$ of gross-substitute valuation functions and a $q$-valid rounding scheme $B$ for the class $\mathcal{C}$, if Mechanism 1 receives valuations from $\mathcal{C}$, then the expected residual surplus of its output is at least
$\frac{q}{r+1}\left(SW(\mathcal{N},\mathcal{I})-\frac{SW(\mathcal{N},2^{r}\cdot\mathcal{I})}{2^{r}}\right).$
@@
We next define the following notation: For an agent $i\in\mathcal{N}$, and a set of items $\mathcal{I}^{\prime}$ which consists of $k$ copies of the original set $\mathcal{I}$, let $p_{i}(\mathcal{N},\mathcal{I}^{\prime})$ be the payment [^5] We require that the payments are consistent with the implementation used in Step 1 of Mechanism 1. charged to agent $i$ in the VCG mechanism when applied to $\mathcal{N}$ and $\mathcal{I}^{\prime}$. We also denote by $p(\mathcal{N},\mathcal{I}^{\prime})=\sum_{i\in\mathcal{N}}p_{i}(\mathcal{N},\mathcal{I}^{\prime})$. Then, we first prove the following claim:
@@
For agents with gross-substitutes valuations it holds that $SW(\mathcal{N},2\cdot\bar{\mathcal{I}})-SW(\mathcal{N},\bar{\mathcal{I}})\geq\frac{p(\mathcal{N},2\cdot\bar{\mathcal{I}})}{2}$.
@@
$\leq\sum_{j\in A_{i}(\mathcal{N},2\cdot\bar{\mathcal{I}})}\left(SW(\mathcal{N},2\cdot\bar{\mathcal{I}}+\{j\})-SW(\mathcal{N},2\cdot\bar{\mathcal{I}})\right),$
@@
$=\frac{q}{r+1}\cdot\left(SW(1)\cdot\left(2-\frac{1}{2^{r}}\right)+\sum_{i=1}^{r}(SW(2^{i})-SW(2^{i-1}))\cdot 2\left(\frac{1}{2^{i}}-\frac{1}{2^{r+1}}\right)\right),$ (3)
where the second equality is by telescoping sum; the third equality is by geometric sum; and the last equality is since $SW(2^{-1})=0$. On the other hand, the expected payment of the mechanism is:
$\mbox{\bf E}\left[\sum_{i\in\mathcal{N}}p_{i}\right]=\frac{q}{r+1}\cdot\sum_{l=0}^{r}\frac{p(2^{l})}{2^{l}}\leq\frac{q}{r+1}\cdot\left(SW(1)+\sum_{l=1}^{r}\frac{SW(2^{l})-SW(2^{l-1})}{2^{l-1}}\right),$ (4)
@@
When Mechanism 1 (implemented with rounding scheme $B_{UD}$) receives unit-demand valuations, the expected residual surplus of its output is at least $O\left(\log(n)\right)$ approximation to the social welfare.
@@
When Mechanism 1 (implemented with rounding scheme $B_{MU}$) receives submodular multi-unit valuations, the expected residual surplus of its output is at least $O\left(\log(n)\right)$ approximation to the social welfare.
We remark that the lower bound of hartline2008optimal immediately implies a lower bound of $\Omega(\log(n))$ for both unit-demand bidders and multi-unit submodular bidders, making the results of Theorem 7 and Theorem 8 asymptotically tight.